OpenStudy (anonymous):

8+x-1/x+4 please show work!

OpenStudy (mertsj):

\[8+\frac{x-1}{x+4}\]

OpenStudy (mertsj):

Is that it? And what are you supposed to do with it?

OpenStudy (anonymous):

yes and add it

OpenStudy (mertsj):

The common denominator is x+4 Multiply 8 by x+4/x+4

OpenStudy (mertsj):

\[8\times\frac{x+4}{x+4}+\frac{x-1}{x+4}\]

OpenStudy (anonymous):

8 X x+-4/x+8?

OpenStudy (mertsj):

\[\frac{8x+32+x-1}{x+4}=\frac{9x+31}{x+4}\]

OpenStudy (anonymous):

thanks